OpenStudy (anonymous):

Solve. -6 + z = 8 z = __

ganeshie8 (ganeshie8):

so you want "z" by itself.

ganeshie8 (ganeshie8):

add "6" both sides

ganeshie8 (ganeshie8):

-6 + z = 8 +6 +6 z = 14